find the nth term of each geometric sequence. a[1]=-10,r=4,n=2
Answer by stanbon(75887)

find the nth term of each geometric sequence. a[1]=-10,r=4,n=2
a(n) = a(1)r^(n-1)
a(2) = -10*4^(2-1) = -10*4 = -40
======================================
Cheers,
Stan H.